36x^2+9y^2+72x-18y-279
how to get standard form

Answer by lwsshak3(11628)

36x^2+9y^2+72x-18y-279
how to get standard form
**
Standard form is an equation, so I assume you forgot to make given expression=0
36x^2+9y^2+72x-18y-279=0
complete the square
36(x^2+2x+1)+9(y^2-2y+1)=279+36+9
36(x+1)^2+9(y-1)^2=324
divide by 324
(x+1)^2/9+(y-1)^2/36=1
This is an equation of an ellipse with vertical major axis with center at (-1,1).
Its standard form: (x-h)^2/b^2+(y-k)^2/a^2=1, a>b, (h,k)=(x,y) coordinates of center.